the perimeter of a rectangle is 58 meters. one side is 3 meters longer than the other side. determine the dimension of the rectangle.

good day! so i tried this equation:
let w be the longer side
let s be the shorter side
2w+2s=58
w=3+s
2(3+s)+s=58
6+s+s=58
6+2s=58
2s=58-6
2s=52
(divide both sides by 2)
s=13
w=13x3
w=39
final answer= 39 and 13

Found 2 solutions by Earlsdon, solver91311:
Answer by Earlsdon(6294)

Your answer is incorrect and your arithmetic is a little strange, to say the least!
w = the shorter side.
s = the longer side.
s = w+3
OK so far! Now:
P = 2(w+s) = 58 Substitute s = w+3 and P = 58.
58 = 2(w+(w+3)) Simplify.
58 = 2(2w+3) Divide both sides by 2.
29 = 2w+3 Subtract 3 from both sides.
26 = 2w Divide by 2.
13 = w and s = 13+3 = 16
Final answer = 16 and 13
Check:
P = 2(w+s) Substitute w = 13 and s = 16.
P = 2(13+16) Simplify.
P = 2(29)
P = 58

Answer by solver91311(24713)



What you have is what some may call a comedy of errors. You have made mistake on top of mistake on top of mistake, and yet you managed to get half of the problem correct.

You said



and



So far, so good. But then you said,



What happened to the factor of 2 in front of the s just before the equals sign?

You should have said



But even if you had that right, you would have been in trouble because you apparently forgot the Distributive Property when you asserted that . Nope. . So, instead of:



you should have said:



Then you made the error that got you back on track. You had and instead of ending up with , you somehow mysteriously reinserted the missing factor of 2 to get , which is what you would have had anyway if you had continued from:





.

But then you went of on a tangent again. Why in Euler, Newton, and Leibnitz's names did you multiply(!?) by 3 instead of adding 3 to like your second original equation said?



BTW:

But

Lesson: ALWAYS check your work. I do.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism